OpenStudy (anonymous):

Math Analysis: Solve the triangle. C=37degrees; a=5.2; b=6

OpenStudy (cwtan):

use cosine rule

OpenStudy (cwtan):

\(a^2=b^2+c^2-2bc \cos A\)

OpenStudy (anonymous):

that formula doesn't look familiar to me

OpenStudy (anonymous):

i know ofc a^2=b^2 + c^2 but not the "-2bc cosA"

OpenStudy (anonymous):

and this is not a right triangle

OpenStudy (anonymous):

r u searching area of triangle

OpenStudy (cwtan):

That's why we use cosine rule

OpenStudy (anonymous):

no, just solving the triangle

OpenStudy (anonymous):

that's the cosine rule?

OpenStudy (cwtan):

ya..... got 3 actually but they all the same for this question \(c^2=a^2+b^2-2ab\cos C\)

OpenStudy (anonymous):

oh, it does sound familiar..

OpenStudy (cwtan):

just substitute the values and you will get the length of c then you can find all others angle and length.

OpenStudy (anonymous):

well if it is not familiar to u then at first just find out the area of triangle by using the very simple formula ie A=1/2 ab sinC then find out the value of c by with another formula ie A= root under s(s-a)(s-b)(s-c) where s is a+b+c/2

OpenStudy (anonymous):

ok, im having a little trouble with finding cosine A

OpenStudy (anonymous):

but cosine law is also one of the formula and may be u'll be familiar with it later

OpenStudy (anonymous):

well there r other formula of cosine laws

OpenStudy (cwtan):

you can also use sine rule \(\Huge\frac{\sin A}{a}=\frac{\sin B}{b}=\frac{\sin C}{c}\)

OpenStudy (anonymous):

like foe this one u use 2abcosC= a^2+b^2-c^2

OpenStudy (cwtan):

or \(\Huge\frac{a}{\sin A}=\frac{b}{\sin B}=\frac{c}{\sin C}\)

OpenStudy (anonymous):

but in order to use that i need to have angle and a side of either A, B, or C right?

OpenStudy (anonymous):

you are given a, b and C so simply use 2abcosC= a^2+b^2-c^2 to find c

OpenStudy (anonymous):

well i got c=3.63

OpenStudy (anonymous):

ok. how about the degrees of A im getting the wrong answer :\

OpenStudy (anonymous):

i got A=60

OpenStudy (anonymous):

it's 59.55 but how did u get that answer